OpenStudy (anonymous):

The sum of the first 30 terms of the sequence an = 6n + 5 is _____.

OpenStudy (anonymous):

i know there is a snap way to do this, but i don't know what it is. i would remember that \[\sum_{k=1}^n k =\frac{n(n-1)}{2}\] and do the following

OpenStudy (a_clan):

a1=11 d=6 n=30

OpenStudy (a_clan):

arithmetic progression

OpenStudy (anonymous):

\[\sum_{k=1}^{30} 6k+5=6\times \sum_{k=1}^{30}k+5\times \sum_{k=1}^{30}1\]

OpenStudy (anonymous):

get \[6\times \frac{30\times 31}{2}+5\times 30\]

OpenStudy (anonymous):

A_clan might have a snappier way

OpenStudy (anonymous):

i get 2940

OpenStudy (amistre64):

\[\frac{n}{2}(first + last)\]

OpenStudy (amistre64):

30(6(1) + 5 + 6(30) + 5) ---------------------- 2 15(6 + 5 + 180 + 5) 15(16 + 180) 15(176) = 2640 is my guess

OpenStudy (amistre64):

to add the numbers 1 to 9 9(1+9) ------ = 90/2 = 45 2 n(first + last) ----------- gets you the sum 2

OpenStudy (amistre64):

lets dbl chk that 4+9+16 = 29 3(4+16) -------- = 3(20)/2 = 60/2 = 30 2 ack! well its close :)